OpenStudy (anonymous):

Identify the correct form of the following sequence. 7 , 28 , 112 , 448 , ...

OpenStudy (anonymous):

u are multiplying by 4 XD

OpenStudy (anonymous):

f(x) = 4x + 7 f(x) = 4(7x) f(x) = 7(4x) f(x) = 4x + 7

OpenStudy (anonymous):

those the answer choices

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know right @***[ISURU]*** it dont look right